OpenStudy (anonymous):

Let f(x)=x^3−2x. Calculate the difference quotient f(2+h)−f(2)/h for h=.1 h=.01 h=−.01 h=−.1 if you can demonstrate how to do one, I will learn to do the rest

OpenStudy (lalaly):

\[f(2+h)=(2+h)^3-2(2+h)\]\[f(2)=2^3-2(2) =8-4=4\]so the first one h=0.1\[f(2+0.1)-\frac{f(2)}{0.1}\]\[=(2+0.1)^3-2(2+0.1)-\frac{4}{0.1}\]and simplify

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The slope of the tangent line to the graph of f(x) at x=2 is m=limh→0 f(2+h)−f(2)/h= ???